The compound phenylthiocarbamide (PTC) tastes very bitter to most people. The

inability to taste PTC is controlled by a single autosomal recessive allele. In the American white

population, about 70% can taste PTC while 30% cannot (are non-tasters). Estimate the

frequencies of the Taster (T) and non-taster (t) alleles in this population. How common are

carriers (those carrying the allele, but not displaying the phenotype) of the non-taster allele?

What assumption did you make to carry out your calculations

how to do i find the carriers?
Answer #1

According to Hardy-Weinberg equation,

p + q = 1

and p2 + q2 + 2pq = 1

where p and q are the frequencies of the two alleles.

Non-taster is the recessive trait, so q2 = 0.3

q = 0.547

p = 1-q = 1 - 0.547 = 0.453

The frequency of carriers = 2pq = 2 * 0.453 * 0.547 = 0.495

Frequency of non-taster = q2 = 0.3

Frequency of taster = p2 + 2pq = 0.7